OpenStudy (anonymous):

You have a bag with three red marbles, two green marbles, and four blue marbles. Suppose you pick one marble at random. What is the probability of picking a blue or green marble?

OpenStudy (anonymous):

3 + 2 + 4 = 9 probability of blue marble : 4 / 9 = 44.4 repeat % of chance probability of green marble : 2/ 9 = 22.2 repeat % of chance